OpenStudy (anonymous):

Write the expression without using parentheses or negative exponents. Assume no variable is zero. (Simplify your answer completely.) (2t^3/t)^-4

OpenStudy (saifoo.khan):

\[(\frac{2t^3}{t})^{-4}\]

OpenStudy (saifoo.khan):

\[(\frac{t}{2t^3})^4\]

OpenStudy (saifoo.khan):

\[\frac{t^4}{2^3t^{3\times 4}}\]

OpenStudy (saifoo.khan):

\[\frac{t^4}{8t^{12}}\]

OpenStudy (saifoo.khan):

\[\frac{1}{8t^8}\]